# Adaptive Multi-Scale Decomposition Framework for Time Series Forecasting

### v1

@@ Abstract
Transformer-based and MLP-based methods have emerged as leading approaches in time series forecasting (TSF). While Transformer-based methods excel in capturing long-range dependencies, they suffer from high computational complexities and tend to overfit. Conversely, MLP-based methods offer computational efficiency and adeptness in modeling temporal dynamics, but they struggle with capturing complex temporal patterns effectively. To address these challenges, we propose a novel MLP-based Adaptive Multi-Scale Decomposition (AMD) framework for TSF. Our framework decomposes time series into distinct temporal patterns at multiple scales, leveraging the Multi-Scale Decomposable Mixing (MDM) block to dissect and aggregate these patterns in a residual manner. Complemented by the Dual Dependency Interaction (DDI) block and the Adaptive Multi-predictor Synthesis (AMS) block, our approach effectively models both temporal and channel dependencies and utilizes autocorrelation to refine multi-scale data integration. Comprehensive experiments demonstrate that our AMD framework not only overcomes the limitations of existing methods but also consistently achieves state-of-the-art performance in both long-term and short-term forecasting tasks across various datasets, showcasing superior efficiency. Code is available at https://github.com/TROUBADOUR000/AMD.
1 Introduction
@@
Time series forecasting (TSF) aims to use historical data to predict future values across various domains, such as finance [1, 2], energy [3], traffic management [4], and weather forecasting [5]. Recently, deep learning has made substantial and reliable advancements in TSF, with the most state-of-the-art performances achieved by Transformer-based methods [6, 7, 8, 9, 10, 11] and MLP-based methods[12, 13, 14, 15, 16].
Transformer-based methods excel at modeling long-range dependencies due to the self-attention mechanisms [17]. Although effective, they come with computational complexity that scales quadratically with the length of the sequence. Additionally, self-attention can diminish the temporal relationships when extracting semantic correlations between pairs in long sequences [13], leading to an overemphasis on mutation points and resulting in overfitting (see Fig. 1(a)). In contrast, MLP-based methods boast significantly lower computational complexity compared to Transformer-based methods. Moreover, MLP-based methods can chronologically model the temporal dynamics in consecutive points, which is crucial for time series analysis [13, 12, 16]. However, the simplicity of linear mappings in existing MLP-based methods presents an information bottleneck [18, 19], hindering their ability to capture diverse temporal patterns and limiting their predictive accuracy.
It is worth noting that time series exhibit distinctly different temporal patterns at various sampling scales [15]. Moreover, the weight of these time scales in predicting future variations is not uniform, as future variations are jointly determined by the entangle of multiple scales (see Fig. 1(a)). For example, weather data sampled hourly reflects fine-grained, sudden changes, while monthly sampled data captures coarse-grained climate variations. Similarly, while short-term hourly data might highlight immediate weather shifts, long-term monthly data provides a broader view of climatic trends. Therefore, efficiently modeling the multi-scale changes in time series and adeptly integrating information across different scales remains a critical challenge.
@@
Figure 1: (a) Illustration of multi-scale temporal patterns in time series and the impact of selector weights. Transformer-based methods often overfit by overemphasizing mutation points, weakening temporal relationships. Efficiently modeling and integrating distinct temporal patterns at various scales is crucial for accurate predictions. (b) Memory usage (MB), training time (ms/iter), and MSE comparisons on the Weather dataset. The input and predicted lengths are set to 512 and 96, respectively. Our proposed AMD achieves a low MSE of 0.145 with 17ms training time and 1349MB memory usage, demonstrating high efficiency and effectiveness.
@@
Motivated by the above observations, we decompose the time series at multiple scales to precisely discern the intertwined temporal patterns within the complex series, rather than merely breaking it down into seasonal and trend components [7, 13]. Subsequently, we model the correlations across different scales in both time and channel dimensions. To account for the varying impacts of different temporal patterns on the future, we employ an autocorrelation approach to model their contributions and adaptively integrate these multi-scale temporal patterns based on their respective influences.
@@
Technically, we propose an MLP-based Adaptive Multi-Scale Decomposition (AMD) Framework to better disentangle and model the diverse temporal patterns within time series. In concrete, the AMD initiates by employing the Multi-Scale Decomposable Mixing (MDM) block, which first decomposes the original time series into multiple temporal patterns through average downsampling and then aggregates these scales to provide aggregate information in a residual way. Subsequently, the Dual Dependency Interaction (DDI) block simultaneously models both temporal and channel dependencies within the aggregated information. Finally, the Adaptive Multi-predictor Synthesis (AMS) block uses the aggregated information to generate specific weights and then employs these weights to adaptively integrate the multiple temporal patterns produced by the DDI. Through comprehensive experimentation, our AMD consistently achieves state-of-the-art performance in both long-term and short-term forecasting tasks, with superior efficiency (see Fig. 1(b)) across multiple datasets.
@@
Our contributions are summarized as follows: (i) We decompose time series across multiple scales to precisely identify intertwined temporal patterns within complex sequences and adaptively aggregate predictions of temporal patterns at different scales, addressing their varied impacts on future forecasts. We also demonstrate the feasibility through theoretical analysis. (ii) We propose a simple but effective MLP-based Adaptive Multi-Scale Decomposition (AMD) framework that initially decomposes time series into diverse temporal patterns, models both temporal and channel dependencies of these patterns, and finally synthesizes the outputs using a weighted aggregation approach to focus on the changes of dominant temporal patterns, thereby enhancing prediction accuracy across scales. (iii) Comprehensive experiments demonstrate that our AMD consistently delivers state-of-the-art performance in both long-term and short-term forecasting across various datasets, with superior efficiency.
2 Related Works
2.1 Time Series Forecasting
@@
Time series forecasting (TSF) involves predicting future values based on past observations in sequential order. In recent years, deep learning methods have gained prominence for their ability to automatically extract intricate patterns and dependencies from data, such as CNN [20, 21, 22], RNN [23, 24], GNN [25, 26, 27], Transformer [7, 6, 10, 28] and MLP [12, 13, 29]. Transformer-based models, renowned for superior performance in handling long and intricate sequential data, have gained popularity in TSF. Autoformer [7] introduces a decomposition architecture and auto-correlation mechanism. PatchTST [6] divides the input time series into patches to enhance the locality of temporal data. In addition to capturing cross-time dependencies, Crossformer [9] also mines cross-variable dependencies to leverage information from associated series in other dimensions. However, Transformer-based models always suffer from efficiency problems due to high computational complexity. In contrast, MLP-based models are more efficient and have a smaller memory footprint. For example, DLinear [13] utilizes the series decomposition as a pre-processing before linear regression and outperforms all of the previous Transformer-based models. FITS [30] proposes a new linear mapping for transformation of complex inputs, with only 10k parameters. However, due to the inherent simplicity and information bottleneck, MLP-based models struggle to effectively capture diverse temporal patterns [19]. In this work, we decompose time series across multiple scales and model the information at each scale with separate linear models, effectively addressing the representational limitations of MLP-based methods.
2.2 Series Decomposition in TSF
@@
Recently, with high sampling rates leading to high-frequency data (such as daily, hourly, or minutely data), real-world time series data often contains multiple underlying temporal patterns. To competently harness different temporal patterns at various scales, several series decomposition designs are proposed [31, 32, 33, 15]. Seasonal Extraction in ARIMA Time Series [34] offers theoretical assurances but is limited to a monthly scale. Seasonal-Trend decomposition [35] is based on moving average and aims to disentangle the seasonal and trend components. FEDformer [8] incorporate frequency information to enhance the series decomposition block. TimesNet [21] decomposes time series data into multiple periods by Fast Fourier Transform, thus several multiple dominant frequencies. SCINet [36] utilizes a hierarchical downsampling tree to iteratively extract and exchange multi-scale information. Following the aforementioned designs, this paper proposes a new Multi-Scale Decomposition method to decompose the time series at multiple scales to discern the intertwined temporal patterns within the complex series precisely.
3 Preliminary: Linear Models with Multi-Scale Information for TSF
@@
After that, we introduce the multi-scale information. For time series forecasting, the most influential real-world applications typically exhibit either smooth or periodicity. Without these characteristics, predictability tends to be low, rendering predictive models unreliable. If the time series only exhibits periodicity, linear models can easily model it [37]. We define the original sequence as $f(x)=f_{0}(x)=[x_{1},x_{2},...,x_{L}]$ and assume that $f(x)$ possesses smoothness. After $k$ downsampling operations with a downsampling rate of $d$, we obtain $n$ sequences $f_{i}(x),\forall i=1,2,...,n$, where $f_{i}(x)=\frac{1}{d}\sum_{j=xd-d+1}^{xd}f_{i-1}(j)$ and $x=1,2,...,\big{[}\frac{L}{d^{i}}\big{]}$. It is noteworthy that $f_{i}(x)\in\mathbb{R}^{C\times\big{[}\frac{L}{d^{i}}\big{]}},\forall i=0,1,...,n$. Then, the sequence $f_{i}(x),\forall i=0,1,...,n$ is transformed into $g_{i}(x)$ through linear mapping and residual calculation. Specifically, $g_{n}(x)=f_{n}(x)$,then through top-down recursion for $i=n-1,...,0$, the operation
@@
is performed recursively, where $W_{i}\in\mathbb{R}^{\big{[}\frac{L}{d^{i+1}}\big{]}\times\big{[}\frac{L}{d^{i}}\big{]}}$. In this case, we derive the Theorem 1 (the proof is available in Appendix A):
@@
This derivation demonstrates that linear models are well-suited to utilize multi-scale temporal patterns in history. For nonperiodic patterns provided they exhibit smoothness, which is often observed in practice, the error remains bounded.
4 Method
As shown in Fig. 2, our proposed AMD mainly consists of three components: Multi-Scale Decomposable Mixing (MDM) Block, Dual Dependency Interaction (DDI) Block, and Adaptive Multi-predictor Synthesis (AMS) Block. Specifically, for the length-$L$ history inputs with $C$ recorded channels, MDM initially decomposes the $\mathbf{X}\in\mathbb{R}^{C\times L}$ into multi-scale temporal patterns through average downsampling and subsequently integrates these patterns residually to yield aggregate information $\mathbf{U}\in\mathbb{R}^{C\times L}$. Next, U is transformed into patches $\hat{\mathbf{U}}\in\mathbb{R}^{C\times N\times P}$, where the length of each patch is $P$ and the number of patches is $N$. The DDI block then takes $\hat{\mathbf{U}}$ as input, concurrently modeling both temporal and channel dependencies, and outputs $\mathbf{V}\in\mathbb{R}^{C\times L}$. Following this, AMS decomposes each channel $\mathbf{u}\in\mathbb{R}^{1\times L}$ of $\mathbf{U}$ from MDM, deriving scores for different temporal patterns and calculating the corresponding weights $\mathbf{S}\in\mathbb{R}^{T\times m}$, where $T$ is the length of the prediction series and $m$ is the number of predictors. It also uses the split outputs $\mathbf{v}\in\mathbb{R}^{1\times L}$ from each channel of $\mathbf{V}$ as inputs for the predictors. These weights, along with the predictor outputs, are then merged through a mixture-of-expert structure to generate the final prediction results. The details of each essential module are explained in the following subsections and the pseudocode is shown in Appendix B.
4.1 Multi-Scale Decomposable Mixing Block
Time series exhibit both coarse-grained temporal patterns and fine-grained patterns. These two types of information influence each other by dynamically interacting: coarse-grained patterns reflect a macroscopic context that shapes fine-grained patterns, while fine-grained patterns offer microscopic feedback that refines and adjusts the coarse-grained patterns [38]. Together, these complementary scales of information provide a comprehensive view of the time series. Therefore, we first decompose the time series into individual temporal patterns, and then mix them to enhance the time series data for a more nuanced analysis and interpretation.
@@ 4.1 Multi-Scale Decomposable Mixing Block
Then, distinct temporal patterns are mixed from the coarse-grained $\boldsymbol{\tau}_{k}$ to the fine-grained $\boldsymbol{\tau}_{1}$ through a feedforward residual network. The mixing of the $i^{th}$ layer of temporal patterns can be represented by the following formula:
$\boldsymbol{\tau}_{i}=\boldsymbol{\tau}_{i}+\text{MLP}(\boldsymbol{\tau}_{i+1})$ (4)
Finally, after completing the mixing of temporal patterns across $h$ scales, we obtain mixed-scale information $\boldsymbol{\tau}_{1}$, with the output of one channel being $\mathbf{u}=\boldsymbol{\tau}_{1}\in\mathbb{R}^{1\times L}$.
4.2 Dual Dependency Interaction Block
Modeling each scale of information separately could ignore the relationships between different scales. In reality, different scales interact with each other, for example in a stock price series where monthly economic trends influence daily fluctuations, and these monthly trends are influenced by annual market cycles.
To model both temporal and channel dependencies between different scales, we propose the DDI block, which first stacks the aggregated information $\mathbf{u}$ from various channels of the MDM into the matrix $\mathbf{U}\in\mathbb{R}^{C\times L}$ and then performs patch operations to transform $\mathbf{U}$ into $\hat{\mathbf{U}}\in\mathbb{R}^{C\times N\times P}$. For each patch, $\hat{\mathbf{V}}_{t}^{t+P}$ represents the embedding output of the residual network, and $\hat{\mathbf{U}}_{t}^{t+P}$ represents a patch of aggregated information from MDM. We adopt a residual block to aggregate historical information to obtain temporal dependency $\mathbf{Z}_{t}^{t+P}$. Next, we perform the transpose operation to fuse cross-channel information through another residual block. Finally, we perform the unpatch operation and split the output information into individual channels to obtain $\mathbf{v}\in\mathbb{R}^{1\times L}$. The interaction of the patch $\hat{\mathbf{U}}_{t}^{t+P}$ can be represented by the following formula:
@@ 4.2 Dual Dependency Interaction Block
where $A^{T}$ is the transpose of matrix A. Finally, by performing the unpatch operation, we obtain the output $\mathbf{V}\in\mathbb{R}^{C\times L}$.
In DDI, dual dependencies are captured under the mixed-scale information $\mathbf{U}$. Temporal dependencies model the interactions across different periods, while cross-channel dependencies model the relationships between different variables, enhancing the representation of the time series. However, through experiments, we find that cross-channel dependencies are not always effective; instead, they often introduce unwanted interference. Therefore, we introduce the scaling rate $\beta$ to suppress the noise.
Figure 2: AMD is an MLP-based model, consisting of the Multi-Scale Decomposable Mixing (MDM) Block, the Dual Dependency Interaction (DDI) Block, and the Adaptive Multi-predictor Synthesis (AMS) Block.
4.3 Adaptive Multi-predictor Synthesis Block
We utilize Mixture-of-Experts (MoE) in AWS due to its adaptive properties, allowing us to design different predictors for different temporal patterns to improve both accuracy and generalization capability. The AMS is partitioned into two components: the temporal pattern selector (TP-Selector) and the temporal pattern projection (TP-Projection). The TP-Selector decomposes different temporal patterns, scores them, and generates the selector weights $\mathbf{S}$. Unlike the downsampling in MDM, which decomposes individual scales to enhance temporal information, the TP-Selector adaptively untangles highly correlated, intertwined mixed scales through feedforward processing. Meanwhile, the TP-Projection synthesizes the multi-predictions and adaptively aggregates the outputs based on the specific weights.
TP-Selector takes a single channel input $\mathbf{u}$ from MDM, decomposes it through feedforward layers, and then applies a noisy gating design [39]:
@@ 4.3 Adaptive Multi-predictor Synthesis Block
$Q(\mathbf{u})=\text{Decompose}(\mathbf{u})+\psi\cdot\text{Softplus}(\text{Decompose}(\mathbf{u})\cdot\mathbf{W}_{\text{noise}})$ (8)

### v2

@@ Abstract
Transformer-based and MLP-based methods have emerged as leading approaches in time series forecasting (TSF). However, real-world time series often show different patterns at different scales, and future changes are shaped by the interplay of these overlapping scales, requiring high-capacity models. While Transformer-based methods excel in capturing long-range dependencies, they suffer from high computational complexities and tend to overfit. Conversely, MLP-based methods offer computational efficiency and adeptness in modeling temporal dynamics, but they struggle with capturing temporal patterns with complex scales effectively. Based on the observation of multi-scale entanglement effect in time series, we propose a novel MLP-based Adaptive Multi-Scale Decomposition (AMD) framework for TSF. Our framework decomposes time series into distinct temporal patterns at multiple scales, leveraging the Multi-Scale Decomposable Mixing (MDM) block to dissect and aggregate these patterns. Complemented by the Dual Dependency Interaction (DDI) block and the Adaptive Multi-predictor Synthesis (AMS) block, our approach effectively models both temporal and channel dependencies and utilizes autocorrelation to refine multi-scale data integration. Comprehensive experiments demonstrate our AMD framework not only overcomes the limitations of existing methods but also consistently achieves state-of-the-art performance across various datasets.
Codehttps://github.com/TROUBADOUR000/AMD
Introduction
@@
Time series forecasting (TSF) aims to use historical data to predict future values across various domains, such as finance (Cheng et al. 2022, 2018; Zhu et al. 2024), energy (Xue and Salim 2023), traffic management (Kim et al. 2022a), and weather forecasting (Volkovs, Urtans, and Caune 2024). Recently, deep learning has made substantial and reliable advancements in TSF, with the most state-of-the-art performances achieved by Transformer-based methods (Nie et al. 2023; Zhou et al. 2022b; Zhang and Yan 2023) and MLP-based methods(Zeng et al. 2023; Tolstikhin et al. 2021; Li et al. 2023b; Qiu et al. 2024).
@@
Figure 1: Illustration of multi-scale temporal patterns in time series and the impact of selector weights. Transformer-based methods often overfit by overemphasizing mutation points, weakening temporal relationships. Efficiently modeling and integrating distinct temporal patterns at various scales is crucial for accurate predictions.
It is worth noting that time series exhibit distinctly different temporal patterns at various sampling scales (Wang et al. 2024). Moreover, the weight of these time scales in predicting future variations is not uniform, as future variations are jointly determined by the entangle of multiple scales (see Fig. 1). For example, weather data sampled hourly reflects fine-grained, sudden changes, while monthly sampled data captures coarse-grained climate variations. Similarly, while short-term hourly data might highlight immediate weather shifts, long-term monthly data provides a broader view of climatic trends. Therefore, efficiently modeling the multi-scale changes in time series and adeptly integrating information across different scales, which requires high-capacity models, remains a critical challenge.
Although Transformer-based methods excel at modeling long-range dependencies due to the self-attention mechanisms (Vaswani et al. 2017), they come with computational complexity that scales quadratically with the length of the sequence. Despite their recognized performance, these methods still face challenges in industrial applications, including low training efficiency and high memory consumption. Additionally, self-attention can diminish the temporal relationships when extracting semantic correlations between pairs in long sequences (Zeng et al. 2023), leading to an overemphasis on mutation points and resulting in overfitting (see Fig. 1). In contrast, MLP-based methods boast significantly lower computational complexity compared to Transformer-based methods. Moreover, MLP-based methods can chronologically model the temporal dynamics in consecutive points, which is crucial for time series analysis (Zeng et al. 2023; Das et al. 2023). However, the simplicity of linear mappings in existing MLP-based methods presents an information bottleneck (Liu et al. 2024b), hindering their ability to capture diverse temporal patterns and limiting their predictive accuracy (Ni et al. 2024).
@@
Motivated by the above observations, we decompose the time series at multiple scales to precisely discern the intertwined temporal patterns within the complex series, rather than merely breaking it down into seasonal and trend components (Wu et al. 2021; Zeng et al. 2023). Subsequently, we model the correlations across different scales in both time and channel dimensions. Unlike average aggregation of TimeMixer (Wang et al. 2024), to account for the varying impacts of different temporal patterns on the future, we employ an autocorrelation approach to model their contributions and adaptively integrate these multi-scale temporal patterns based on their respective influences.
@@
Technically, we propose an MLP-based Adaptive Multi-Scale Decomposition (AMD) Framework to better disentangle and model the diverse temporal patterns within time series. In concrete, the AMD initiates by employing the Multi-Scale Decomposable Mixing (MDM) block, which first decomposes the original time series into multiple temporal patterns through average downsampling and then aggregates these scales to provide aggregate information in a residual way. Subsequently, the Dual Dependency Interaction (DDI) block simultaneously models both temporal and channel dependencies within the aggregated information. Finally, the Adaptive Multi-predictor Synthesis (AMS) block uses the aggregated information to generate specific weights and then employs these weights to adaptively integrate the multiple temporal patterns produced by the DDI. Through comprehensive experimentation, our AMD consistently achieves state-of-the-art performance in both long-term and short-term forecasting tasks, with superior efficiency across multiple datasets.
Our contributions are summarized as follows:
  • We decompose time series across multiple scales to precisely identify dominant intertwined temporal patterns within complex sequences and adaptively aggregate predictions of temporal patterns at different scales, addressing their varied impacts on future forecasts. We also demonstrate the feasibility through theoretical analysis.
@@
  • We propose a simple but effective MLP-based Adaptive Multi-Scale Decomposition (AMD) framework that initially decomposes time series into diverse temporal patterns, models both temporal and channel dependencies of these patterns, and finally synthesizes the outputs using a weighted aggregation approach to focus on the changes of dominant temporal patterns, thereby enhancing prediction accuracy across scales.
  • Comprehensive experiments demonstrate that our AMD consistently delivers state-of-the-art performance in both long-term and short-term forecasting across various datasets, with superior efficiency.
Related Works
Time Series Forecasting
@@
Recently, deep learning methods for TSF have gained prominence , such as CNN (Wang et al. 2023; Dai et al. 2024), RNN (Lin et al. 2023; Jia et al. 2023), GNN (Wu et al. 2020; Yi et al. 2023), Transformer (Wu et al. 2021; Liu et al. 2024a) and MLP (Das et al. 2023; Zhang et al. 2022). Transformer-based models, renowned for superior performance in handling long sequential data, have gained popularity in TSF. PatchTST (Nie et al. 2023) divides the time series into patches to enhance locality. Beyond cross-time dependencies, Crossformer (Zhang and Yan 2023) also mines cross-variable dependencies. However, Transformer-based models always suffer from efficiency problems due to high computational complexity. In contrast, MLP-based models have a smaller memory footprint. FITS (Xu, Zeng, and Xu 2024) proposes a new linear mapping for the transformation of complex inputs, with only 10k parameters. However, due to the inherent simplicity and information bottleneck, MLP-based models struggle to effectively capture diverse temporal patterns (Ni et al. 2024). In this work, we decompose time series across multiple scales and use separate linear models for each, effectively addressing such representational limits.
Mixture of Experts for TSF
The concept of Mixture of Experts (MoE) has a long history (Jacobs et al. 1991; Jordan 1994). Recent advances in large language models (LLM) have reignited interest in MoE (Kim, Lim, and Han 2024; Xue et al. 2024). However, in the TFS domain, the utilization of MoE remains limited. FiLM (Zhou et al. 2022a) assigns input sequences with different time horizons to various experts for forecasting. FEDformer (Zhou et al. 2022b) extracts a set of data-dependent weights by MoE for combining multiple trend components as the final trend. In our work, we harness the adaptability of MoE to craft distinct predictors customized for individual temporal patterns, thereby optimizing model performance and enhancing overall model interpretability.
Series Decomposition in TSF
@@
Lately, with high sampling rates leading to high-frequency data (such as daily, hourly, or minutely data), real-world time series data often contains multiple underlying temporal patterns. To competently harness different temporal patterns at various scales, several series decomposition designs are proposed (Liu et al. 2022; Oreshkin et al. 2020). Seasonal-Trend decomposition (Wu et al. 2021) uses moving averages to separate seasonal and trend components. TimesNet (Wu et al. 2023) uses Fast Fourier Transform to extract multiple dominant frequencies, while SCINet (LIU et al. 2022) employs a hierarchical downsampling tree to iteratively decompose multi-scale information. TimeMixer (Wang et al. 2024) downsamples time series into different scales of seasonal and trend sequences. However, they ignore high-order interactions across different scales. Refining these designs, we introduces an adaptive multi-scale decomposition method to model the dynamic interactions and accurately discern dominated temporal patterns.
Figure 2: Overall structure of the proposed AMD. MDM decomposes and mixes multi-scale information. DDI models dynamic interactions across different scales. AMS discerns dominated temporal patterns and adaptively makes predictions.
Preliminary: Linear Models with Multi-Scale Information for TSF
@@
After that, we introduce the multi-scale information. For time series forecasting, the most influential real-world applications typically exhibit either smooth or periodicity. Without these characteristics, predictability tends to be low, rendering predictive models unreliable. If the time series only exhibits periodicity, linear models can easily model it (Chen et al. 2023). We define the original sequence as $f(x)=f_{0}(x)=[x_{1},x_{2},...,x_{L}]$ and assume that $f(x)$ possesses smoothness. After $k$ downsampling operations with a downsampling rate of $d$, we obtain $n$ sequences $f_{i}(x),\forall i=1,2,...,n$, where $f_{i}(x)=\frac{1}{d}\sum_{j=xd-d+1}^{xd}f_{i-1}(j)$ and $x=1,2,...,\big{[}\frac{L}{d^{i}}\big{]}$. It is noteworthy that $f_{i}(x)\in\mathbb{R}^{C\times\big{[}\frac{L}{d^{i}}\big{]}},\forall i=0,1,...,n$. Then, the sequence $f_{i}(x),\forall i=0,1,...,n$ is transformed into $g_{i}(x)$ through linear mapping and residual calculation. Specifically, $g_{n}(x)=f_{n}(x)$,then through top-down recursion for $i=n-1,...,0$, the operation
@@
is performed recursively, where $W_{i}\in\mathbb{R}^{\big{[}\frac{L}{d^{i+1}}\big{]}\times\big{[}\frac{L}{d^{i}}\big{]}}$. In this case, we derive the Theorem 1
@@
This derivation demonstrates that linear models are well-suited to utilize multi-scale information. For nonperiodic patterns provided they exhibit smoothness, which is often observed in practice, the error remains bounded.
Method
As mentioned above, the key challenge for accurate forecasting is to deal with multi-scale temporal patterns that have an entangled influence on the future evolution of variables. In this paper, we propose AMD to enhance forecasting accuracy by capturing multi-scale entanglement through multi-scale decomposition and an autocorrelation approach (see Fig. 2). Specifically, AMD mainly consists of three components: Multi-Scale Decomposable Mixing (MDM) Block, Dual Dependency Interaction (DDI) Block, and Adaptive Multi-predictor Synthesis (AMS) Block. The details of each essential module are explained in the following subsections.
Multi-Scale Decomposable Mixing Block
Time series exhibit both coarse-grained temporal patterns and fine-grained patterns. Together, these complementary scales of information provide a comprehensive view of the time series. Therefore, we first decompose the time series into individual temporal patterns, and then mix them to enhance the time series data for a more nuanced analysis and interpretation.
@@
Then, distinct temporal patterns are mixed from the coarse-grained $\boldsymbol{\tau}_{h}$ to the fine-grained $\boldsymbol{\tau}_{1}$ through a feedforward residual network, while $\xi_{i}$ represents the mixed data and $\xi_{h}=\boldsymbol{\tau}_{h}$. The mixing of the $i^{th}$ layer of temporal patterns can be represented by the following formula:
$\boldsymbol{\xi}_{i}=\boldsymbol{\tau}_{i}+\text{MLP}(\boldsymbol{\xi}_{i+1})$ (4)
Finally, after completing the mixing of temporal patterns across $h$ scales, we obtain mixed-scale information $\boldsymbol{\xi}_{1}$, with the output of one channel being $\mathbf{u}=\boldsymbol{\xi}_{1}\in\mathbb{R}^{1\times L}$.
Dual Dependency Interaction Block
We observed potential dynamic interactions across different scales in various time series. Intuitively, in a stock price series, daily fluctuations are influenced by monthly economic trends, which in turn are shaped by annual market cycles. As a result, changes at each scale propagate through the system, with the effects at one scale spilling over and altering the dynamics at other scales. However, ScaleFormer (Shabani et al. 2023) and TimeMixer (Wang et al. 2024) ignore these potential high-order interactions among different scales.
To model the interactions across different scales, including both temporal and channel dependencies, we propose the DDI block with time-mixing and channel-mixing. DDI first stacks the aggregated information $\mathbf{u}$ from various channels of the MDM into the matrix $\mathbf{U}\in\mathbb{R}^{C\times L}$ and then performs patch operations to transform $\mathbf{U}$ into $\hat{\mathbf{U}}\in\mathbb{R}^{C\times N\times P}$. For each patch, $\hat{\mathbf{V}}_{t}^{t+P}$ represents the embedding output of the residual network, and $\hat{\mathbf{U}}_{t}^{t+P}$ represents a patch of aggregated information from MDM. We adopt an MLP shared by time steps to aggregate time-mixing information through the time dimension for each channel to obtain temporal dependencies $\mathbf{Z}_{t}^{t+P}$. Next, we perform the transpose operation to fuse channel-mixing information along the time domain through another MLP shared by channels. Finally, we perform the unpatch operation and split the output information into individual channels to obtain $\mathbf{v}\in\mathbb{R}^{1\times L}$. The residual operation ensures that the model retains its ability to capture temporal dependencies while effectively exploiting cross-channel dependencies. The interaction of the patch $\hat{\mathbf{U}}_{t}^{t+P}$ can be represented by the following formula:
@@
where $A^{T}$ is the transpose of matrix A.
In DDI, dual dependencies are captured under the mixed-scale information $\mathbf{U}$. Temporal dependencies model the interactions across different periods, while cross-channel dependencies model the relationships between different variables. However, through experiments, we find that cross-channel dependencies are not always effective, especially when the target time series is not correlated with other covariates; instead, they often introduce unwanted interference. Therefore, we introduce the scaling rate $\beta$ to suppress the noise and balance the emphasis on temporal dependencies and cross-channel dependencies.
Adaptive Multi-predictor Synthesis Block
It is noted that the dominant temporal patterns, which significantly influence future variables, dynamically change over different periods and have stronger predictive power. In contrast to the average aggregation in TimeMixer (Wang et al. 2024), which overlooks these dominant changes, we exploit the adaptive properties of MoE to design specific predictors for each temporal pattern. By dynamically assigning more attention to the dominant scales, we improve both accuracy and generalisability.
The AMS is partitioned into two components: the temporal pattern selector (TP-Selector) and the temporal pattern projection (TP-Projection). The TP-Selector decomposes different temporal patterns and generates the selector weights $\mathbf{S}$. Unlike the downsampling in MDM, which decomposes individual scales to enhance temporal information, the TP-Selector adaptively untangles highly correlated, intertwined mixed scales through feedforward processing. Meanwhile, the TP-Projection synthesizes the multi-predictions and adaptively aggregates the outputs based on the specific weights.
TP-Selector takes a single channel input $\mathbf{u}$ from MDM, decomposes it through feedforward layers, and then applies a noisy gating design (Shazeer et al. 2017):
@@
$Q(\mathbf{u})=\text{Decomp.}(\mathbf{u})+\psi\cdot\text{Softplus}(\text{Decomp.}(\mathbf{u})\cdot\mathbf{W}_{\text{noise}})$ (8)